please help me solve this equation, i dont get how to do it
2(2x+3)=6

Answer by Here2Help(19)

Here you go:
Original problem--
2(2x + 3) = 6
Reorder the terms:
2(3 + 2x) = 6
(3 * 2 + 2x * 2) = 6
(6 + 4x) = 6
Add '-6' to each side of the equation.
6 + -6 + 4x = 6 + -6
Combine like terms: 6 + -6 = 0
0 + 4x = 6 + -6
4x = 6 + -6
Combine like terms: 6 + -6 = 0
4x = 0
Solving
4x = 0
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Divide each side by '4'.
x = 0
Final answer:
x = 0